Hart Labs, Inc. provides mad cow disease testing for both state and federal governmental agricultural agencies. Because the cV Your answer is correct. Compute the total overhead variance. Favorable Total Overhead variance 480Prepare an income statement for management. HART LABS, INC. Income Statement For the Month Ended November 30, 2020 V Sales ReHart Labs, Inc. provides mad cow disease testing for both state and federal governmental agricultural agencies. Because the c
Answer #1
Explanation for the Unfavourable Variances:
1..
For D/M Quantity Variance
Test tubes consumed /test has exceeded the standard by (1680/800)=2.1-2= 0.1 per test tube--resulting in a total UNF- variance due to quantity of 800*0.1* $ 1=80
For D/L Quantity Variance
Labor hrs. consumed /test has exceeded the standard by (824/800)=1.03-1= 0.03 hrs. per test --resulting in a total UNF- variance due to labor hrs. of 800*0.03*30= 720
2..
For D/M Quantity Variance
Test tubes consumed /test has exceeded the standard by (3050/1475)=2.068-2= 0.068 per test tube--resulting in a total UNF- variance due to quantity of 1475*0.068*1.46= 146
For D/L Quantity Variance
Labor hrs. consumed /test has exceeded the standard by (1550/1475)=1.0508-1= 0.0508 hrs. per test --resulting in a total UNF- variance due to labor hrs. of 1475*0.0508*24= 1800
Net OH variance (F) for 2..
Std./bud. Actuals F/(UF)
VOH 8850 7400 1450
FOH 14000 15000 -1000
Total Ohs 22850 22400 450
2.. Hart Labs inc.
Income statement
Sales revenue 75000
Cost of service provided(1475*42.92) 63307
Gross profit(at std.) 11693
Variances:
MPV 183
LPV 1550
OH V 450
MQV -146
LQV -1800
Total variance 237
Gross profit(actual) 11930
Sell.& admn. Exp. 5000
Net income/(loss) 6930